# Energy-Efficient Designs for SIM-Based Broadcast MIMO Systems

### v1

@@ IV-A Precoding Matrix Optimization
where $\mathbf{Z}_{k}=\bar{\mathbf{H}}_{k}\mathbf{X}_{k}\in\mathbb{C}^{N_{r}\times N_{r}}$, $\mathbf{\hat{Z}}_{k}=\bar{\mathbf{H}}_{k}\mathbf{X}_{k}^{(n)}\in\mathbb{C}^{N_{r}\times N_{r}}$, $\mathbf{Y}_{k}=\sum\nolimits_{j=1,j\neq k}^{K}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}\mathbf{X}_{j}^{H}\bar{\mathbf{H}}_{k}^{H}+\mathbf{I}\in\mathbb{C}^{N_{r}\times N_{r}}$, $\hat{\mathbf{Y}}_{k}=\sum\nolimits_{j=1,j\neq k}^{K}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}^{(n)}(\mathbf{X}_{j}^{(n)})^{H}\bar{\mathbf{H}}_{k}^{H}+\mathbf{I}\in\mathbb{C}^{N_{r}\times N_{r}}$ and $\mathbf{A}_{k}=\hat{\mathbf{Y}}_{k}^{-1}-(\hat{\mathbf{Z}}_{k}\hat{\mathbf{Z}}_{k}^{H}+\hat{\mathbf{Y}}_{k})^{-1}\in\mathbb{C}^{N_{r}\times N_{r}}$. Consequently, the resulting approximate problem of (44) is given by
@@
Implementing the Karush-Kuhn-Tucker (KKT) first-order optimality condition to (48) by taking the gradient of this expression w.r.t. $\mathbf{X}_{j}^{*}$ and setting it to 0, we obtain
@@ 14          $n\leftarrow n+1$
16           until $f(\mathbf{X}^{(n+1)})-f(\mathbf{X}^{(n)})>\epsilon_{1,L}$
@@ 21          $\mathbf{X}^{(0)}\leftarrow\mathbf{X}^{(n)}$
23           until  $\lambda_{L}^{(m+1)}-\lambda_{L}^{(m)}>\epsilon_{2,L}$
@@ 35                         $\mathbf{P}_{k}=\mathbf{H}^{H}\mathbf{X}_{\text{opt},k}$
To solve for $\mathbf{X}_{j}$ we differentiate two cases. If $\mathbf{H}$ is row-rank matrix, e.g. (i.e., $KN_{r}<N_{t}$), then $\bar{\mathbf{H}}$ is invertible, and thus (50) results in
@@
$\mathbf{H}\Bigl{(}\mathbf{H}_{j}^{H}\hat{\mathbf{Y}}_{j}^{-1}\hat{\mathbf{Z}}_{j}-\sum_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{A}_{k}^{H}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}-\lambda_{L}\mathbf{H}^{H}\mathbf{X}_{j}\Bigr{)}=\boldsymbol{0}$ (52)
@@
$\mathbf{X}_{j}=\Bigl{(}\sum_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{A}_{k}^{H}\bar{\mathbf{H}}_{k}+\lambda_{L}\mathbf{H}^{H}\Bigr{)}^{+}\mathbf{H}_{j}^{H}\hat{\mathbf{Y}}_{j}^{-1}\hat{\mathbf{Z}}_{j}.$ (53)
If the obtained solution from (51) or (53) satisfy the power constraint (44b) then it also the general case solution. Otherwise, the optimal $\mathbf{X}$ is the solution of the achievable rate optimization problem
$\underset{\mathbf{X}}{\operatorname{maximize}}$ $\sum_{k=1}^{K}\bar{R}_{k}$ (54a)
$\operatorname{subject~{}to}$ $\sum_{k=1}^{K}\operatorname{Tr}(\bar{\mathbf{H}}\mathbf{X}_{k}\mathbf{X}_{k}^{H})\leq P_{\max},$ (54b)
which can be solved by [26, Algorithm 1]. The described optimization algorithm is summarized in Algorithm 3.
@@ IV-B SIM Phase Shift Optimization
For fixed $\mathbf{P}$, the SIM phase shift optimization problem is formulated as
$\underset{\boldsymbol{\phi}}{\operatorname{maximize}}\quad$ $\tau(\boldsymbol{\phi})=\sum\nolimits_{k=1}^{K}R_{k}(\boldsymbol{\phi})$ (55a)
$\operatorname{subject~{}to}\quad$ $\left|\boldsymbol{\phi}\right|=1$ (55b)
where the achievable rate for user $k$ is expressed as
$R_{\text{L},k}(\boldsymbol{\phi})$ $=\ln\left|\mathbf{I}+\sum\nolimits_{j=1}^{K}\mathbf{H}_{k}\mathbf{P}_{j}\mathbf{P}_{j}^{H}\mathbf{H}_{k}^{H}\right|$
$-\ln\left|\mathbf{I}+\sum\nolimits_{j=1,j\neq k}^{K}\mathbf{H}_{k}\mathbf{P}_{j}\mathbf{P}_{j}^{H}\mathbf{H}_{k}^{H}\right|.$ (56)
Since we again apply the projected gradient method for the phase shift optimization, the rate expression in (9) is used, instead of (45). The obvious reason is that it is easier to find the gradient of the objective w.r.t $\boldsymbol{\phi}$ using (9) than using (9).
The optimization of the phase shifts of the BS SIM follows the same steps as in the case of DPC in subsection III-B. Specifically, the phase shifts are iteratively updated as
Theorem 2.
$\nabla_{\boldsymbol{\phi}^{l}}\tau(\boldsymbol{\phi})=\operatorname{vec}_{d}\left(\sum_{k=1}^{K}\boldsymbol{\Theta}^{l+1:L}\mathbf{G}_{k}^{H}\mathbf{F}_{1,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{s}\boldsymbol{\Theta}^{1:l-1}\mathbf{W}_{l}^{H}\right)$
$-\operatorname{vec}_{d}\left(\sum_{k=1}^{K}\boldsymbol{\Theta}^{l+1:L}\mathbf{G}_{k}^{H}\mathbf{F}_{2,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{k}\boldsymbol{\Theta}^{1:l-1}\mathbf{W}_{l}^{H}\right)$ (58)
Proof:
See Appendix A. ∎
The outline of the proposed algorithm is given in Algorithm 4.
@@ 4     Call Algorithm 3 to obtain $\mathbf{P}^{(n+1)}$
10          if  $\tau(\boldsymbol{\phi}^{(n+1)})<\tau(\boldsymbol{\phi}^{(n)})+\delta||\boldsymbol{\phi}^{(n+1)}-\boldsymbol{\phi}^{(n)}||^{2}$  then
@@ 12               $t_{n}\leftarrow\rho t_{n}$
16                until  $\tau(\boldsymbol{\phi}^{(n+1)})\geq\tau(\boldsymbol{\phi}^{(n)})+\delta||\boldsymbol{\phi}^{(n+1)}-\boldsymbol{\phi}^{(n)}||^{2}$
@@ 17               $n\leftarrow n+1$
19                until convergence of $\eta$ in (13a)
Algorithm 4 Proposed algorithm for the EE optimization for a SIM-aided broadcast system with linear precoding.
@@ V Computational Complexity
In this section, the computational complexity for SIM-aided broadcast systems with DPC and linear precoding are obtained by counting the required number of complex multiplications. In the following complexity analysis, for ease of exposition, we assume that $N,N_{t}\gg N_{r}$ which is the typical case for a SIM-based broadcast communication system.
@@ V-A DPC Precoding
The optimization of the covariance matrices is performed by Algorithm 1. The complexity of obtaining $\mathbf{U}$ from $\mathbf{S}$ can be neglected. In addition, the complexity of calculating all $\hat{\mathbf{V}}_{j}^{(n)}$, $\hat{\mathbf{Y}}_{j}^{(n)}$, $\mathbf{A}_{j}^{(n)}$ and $\mathbf{B}_{j}^{(n)}$ is $\mathcal{O}(KN_{t}N_{r}^{2})$ as explained previously. Furthermore, $\mathcal{O}(KN_{t}^{2}N_{r})$ multiplications is needed to calculate all $\mathbf{U}_{j}$ in (32). The complexity of calculating the objective function is $\mathcal{O}(KN_{t}^{2}N_{r}+N_{t}^{3})$. Let $I_{U}$ be the number of inner loops (i.e., lines 3 to 7 in Algorithm 1). Then the complexity of lines 1 to 11 in Algorithm 1 is $\mathcal{O}(I_{U}(KN_{t}^{2}N_{r}+N_{t}^{3}))$.
For optimizing the SIM phase shifts, we need $\mathcal{O}(KNN_{t}N_{r})$ multiplications for the computation of $\sum\nolimits_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k}$. The complexity of the matrix inversion $\mathbf{A}^{-1}$ and its multiplication with the previous sum is $\mathcal{O}(N_{t}^{3}+NN_{t}^{2})$. As all the matrix product $(\mathbf{W}^{m})^{H}(\boldsymbol{\mathbf{\Phi}}^{m})^{H}$ are precomputed in advance and the fact that we need only the diagonal elements in (37), the additional complexity is $\mathcal{O}(LN^{3})$. Hence, the complexity of the gradient calculation is $\mathcal{O}(NN_{t}^{2}+LN^{3})$. After obtaining $\boldsymbol{\phi}^{(n+1)}$, the complexity of calculating $\mathbf{B}$ and $\mathbf{H}$ is $\mathcal{O}(LN^{3})$. In addition, $\mathcal{O}(KN_{t}N_{r}(N_{t}+N_{r})+N_{t}^{3})\approx\mathcal{O}(N_{t}^{3})$ multiplications is needed for calculating $h(\boldsymbol{\phi}^{(n+1)})$. The complexity of optimizing the SIM phase shifts is given by $\mathcal{O}(NN_{t}^{2}+LN^{3}+I_{\phi,D}(LN^{3}+N_{t}^{3}))\approx\mathcal{O}(NN_{t}^{2}+I_{\phi,D}(LN^{3}+N_{t}^{3}))$, where $I_{\phi,D}$ is the number of line search loops.
Therefore, the overall computational complexity for one iteration of the DPC algorithm is given by
$C_{\text{DPC}}$ $=\mathcal{O}(I_{U}(KN_{t}^{2}N_{r}+N_{t}^{3})+NN_{t}^{2}$
$+I_{\phi,D}(LN^{3}+N_{t}^{3})).$ (59)
V-B Linear Precoding
The optimization of the precoding matrices is specified by Algorithm 3. The calculation of $\bar{\mathbf{H}}$ requires $\mathcal{O}(K^{2}N_{t}N_{r}^{2})$ multiplications. Computing the initial $\mathbf{X}$ from $\mathbf{P}$ requires $\mathcal{O}(K^{2}N_{t}N_{r}^{2}+K^{3}N_{r}^{3})$ multiplications. Furthermore, the complexity of calculating all $\mathbf{\hat{Z}}_{k}$, $\mathbf{\hat{Y}}_{k}$ and $\mathbf{A}_{k}$ is $\mathcal{O}(K^{3}N_{r}^{3}+K^{2}N_{r}^{3})$. Next, we need to determine the optimal $\mathbf{X}$ according to (51) if $N_{t}\geq KN_{r}$, or otherwise according to (53). The calculation of the optimal $\mathbf{X}$ according to (51) requires $\mathcal{O}(K^{3}N_{r}^{3})$ multiplications and according to (53) $\mathcal{O}(K^{3}N_{r}^{3}+K^{3}N_{t}N_{r}^{2})\approx\mathcal{O}(K^{3}N_{t}N_{r}^{2})$, which can be written as
@@ V-B Linear Precoding
The complexity for calculating $\sum_{k}\bar{R}_{k}$ and $f(\mathbf{X}^{(n)})$ can be neglected. If the number of inner loops (i.e., lines 3 to 8 in Algorithm 3) is $I_{X}$, then the complexity of Algorithm 3 is given by $\mathcal{O}(I_{X}C_{x})$.
To optimize the phase shifts of the SIM, we need $\mathcal{O}(KN_{t}^{2})$ multiplications to obtain $\hat{\mathbf{P}}_{s}$ and all $\hat{\mathbf{P}}_{k}$ matrices. The complexity of calculating $\mathbf{F}_{1,k}$ and $\mathbf{F}_{2,k}$ is $\mathcal{O}(KN_{t}N_{r}(N_{t}+N_{r}))$ and the same is also true for $\mathbf{F}_{1,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{s}-\mathbf{F}_{2,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{k}$. Multiplying these terms with $\mathbf{G}_{k}^{H}$ has the complexity of $\mathcal{O}(KNN_{t}^{2})$. Utilizing the fact that all the matrix product $(\mathbf{W}^{m})^{H}(\boldsymbol{\mathbf{\Phi}}^{m})^{H}$ are precomputed and that we need only the diagonal elements in (58), the additional complexity is $\mathcal{O}(LN^{3})$. Hence, the complexity of the gradient calculation is $\mathcal{O}(KNN_{t}^{2}+LN^{3})$. After obtaining $\boldsymbol{\phi}^{(n+1)}$, the complexity of calculating $\mathbf{B}$ and $\mathbf{H}$ is $\mathcal{O}(LN^{3})$. To obtain all terms $\sum\nolimits_{j}\mathbf{H}_{k}\mathbf{P}_{j}\mathbf{P}_{j}^{H}\mathbf{H}_{k}^{H},$ we need $\mathcal{O}(KN_{t}N_{r}(N_{t}+N_{r}))$ multiplications. Any additional complexity for computing $g(\boldsymbol{\phi}^{(n+1)})$ can be neglected. Hence, the complexity of optimizing the SIM phase shifts is given by $\mathcal{O}(KNN_{t}^{2}+LN^{3}+I_{\phi,L}LN^{3})\approx\mathcal{O}(KNN_{t}^{2}+I_{\phi,L}LN^{3})$, where $I_{\phi,L}$ is the number of line search loops.
Therefore, the overall computational complexity for one iteration of the linear precoding algorithm is given by
$C_{\text{LIN}}=\mathcal{O}(I_{X}C_{x}+KNN_{t}^{2}+I_{\phi,L}LN^{3}).$ (61)

### v2

@@ IV-A2 SCA Method: A Tight Lower Bound of Sum Rate
where $\mathbf{Z}_{k}=\bar{\mathbf{H}}_{k}\mathbf{X}_{k}\in\mathbb{C}^{N_{r}\times N_{r}}$, $\mathbf{\hat{Z}}_{k}=\bar{\mathbf{H}}_{k}\mathbf{X}_{k}^{(n)}\in\mathbb{C}^{N_{r}\times N_{r}}$, $\mathbf{Y}_{k}=\sum\nolimits_{j=1,j\neq k}^{K}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}\mathbf{X}_{j}^{H}\bar{\mathbf{H}}_{k}^{H}+\mathbf{I}\in\mathbb{C}^{N_{r}\times N_{r}}$, $\hat{\mathbf{Y}}_{k}=\sum\nolimits_{j=1,j\neq k}^{K}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}^{(n)}(\mathbf{X}_{j}^{(n)})^{H}\bar{\mathbf{H}}_{k}^{H}+\mathbf{I}\in\mathbb{C}^{N_{r}\times N_{r}}$, $\mathbf{A}_{k}=\hat{\mathbf{Y}}_{k}^{-1}-(\hat{\mathbf{Z}}_{k}\hat{\mathbf{Z}}_{k}^{H}+\hat{\mathbf{Y}}_{k})^{-1}\in\mathbb{C}^{N_{r}\times N_{r}}$, and $\mathbf{X}_{k}^{(n)}$ represents the value of $\mathbf{X}_{k}$ at the $n$-th iteration. Consequently, the resulting approximate problem of (42) is given by
@@ IV-A3 Dinkelbach’s method: Closed-form Solution
Applying the Karush-Kuhn-Tucker (KKT) first-order optimality condition to (46), we take the gradient w.r.t. $\mathbf{X}_{j}^{*}$ and set it to 0, giving
@@ 14    $n\leftarrow n+1$
16     until $\lambda_{L}^{(m+1)}-\lambda_{L}^{(m)}>\epsilon_{2,L}$
@@ 21    $\mathbf{X}^{(0)}\leftarrow\mathbf{X}^{(n)}$
23     until $f(\mathbf{X}^{(n+1)})-f(\mathbf{X}^{(n)})>\epsilon_{1,L}$
@@ 35          $\mathbf{P}_{k}=\mathbf{H}^{H}\mathbf{X}_{\text{opt},k}$
To solve for $\mathbf{X}_{j}$ we differentiate two cases. If $\mathbf{H}$ is a row-rank matrix, e.g. (i.e., $KN_{r}<N_{t}$), then $\bar{\mathbf{H}}$ is invertible, and thus (48) results in
@@
$\mathbf{H}\Bigl(\mathbf{H}_{j}^{H}\hat{\mathbf{Y}}_{j}^{-1}\hat{\mathbf{Z}}_{j}-\sum\nolimits_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{A}_{k}^{H}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}-\lambda_{L}\mathbf{H}^{H}\mathbf{X}_{j}\Bigr)=\boldsymbol{0}$ (50)
@@
$\mathbf{X}_{j}=\Bigl(\sum\nolimits_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{A}_{k}^{H}\bar{\mathbf{H}}_{k}+\lambda_{L}\mathbf{H}^{H}\Bigr)^{+}\mathbf{H}_{j}^{H}\hat{\mathbf{Y}}_{j}^{-1}\hat{\mathbf{Z}}_{j}.$ (51)
If the solution obtained from (49) or (51) satisfies the power constraint (42b), then it is also the optimal solution to (42). Otherwise, the optimal $\mathbf{X}$ is determined by solving the achievable sum rate optimization problem
$\underset{\mathbf{X}}{\operatorname{maximize}}$ $\sum\nolimits_{k=1}^{K}\bar{R}_{k}$ (52a)
$\operatorname{subject~to}$ $\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\bar{\mathbf{H}}\mathbf{X}_{k}\mathbf{X}_{k}^{H})\leq P_{\max},$ (52b)
which can be solved by [24, Algorithm 1]. The described optimization algorithm is summarized in Algorithm 3.
@@ IV-B SIM Phase Shift Optimization
For a fixed $\mathbf{P}$, the SIM phase shift optimization problem is formulated as
$\underset{\boldsymbol{\phi}}{\operatorname{maximize}}\quad$ $\tau(\boldsymbol{\phi})=\sum\nolimits_{k=1}^{K}R_{\text{L},k}(\boldsymbol{\phi})$ (53a)
$\operatorname{subject~to}\quad$ $\left|\phi_{n}^{l}\right|=1,\forall l,n,$ (53b)
where we have used the rate expression in (8) instead of (43). This choice not only simplifies the gradient computation of the objective function w.r.t $\boldsymbol{\phi}$ but also leads to faster convergence rate as explained in [25].
The optimization of the phase shifts of the BS SIM follows the same steps as in the case of DPC discussed in subsection III-B. Specifically, the phase shifts are iteratively updated according to
@@ 4  Call Algorithm 3 to obtain $\mathbf{P}^{(n+1)}$
[…]
@@ 12      $t_{n}\leftarrow\rho t_{n}$
[…]
@@ 17      $n\leftarrow n+1$
19       until convergence of $\eta_{\mathrm{lp}}$ in (12a)
Algorithm 4 Proposed algorithm for the EE optimization for a SIM-aided broadcast system with LP.
Theorem 2.
$\nabla_{\boldsymbol{\phi}^{l}}\tau(\boldsymbol{\phi})=\operatorname{vec}_{d}\left(\sum\nolimits_{k=1}^{K}\boldsymbol{\Theta}^{l+1:L}\mathbf{G}_{k}^{H}\mathbf{F}_{1,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{s}\boldsymbol{\Theta}^{1:l-1}\mathbf{W}_{l}^{H}\right)$
$-\operatorname{vec}_{d}\left({}_{k=1}^{K}\sum\nolimits_{k=1}^{K}\boldsymbol{\Theta}^{l+1:L}\mathbf{G}_{k}^{H}\mathbf{F}_{2,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{k}\boldsymbol{\Theta}^{1:l-1}\mathbf{W}_{l}^{H}\right)$ (55)
Proof:
See the Appendix. ∎
The outline of the proposed algorithm is given in Algorithm 4.
@@ V Computational Complexity
In this section, the computational complexity for SIM-aided broadcast systems with DPC and LP is obtained by counting the required number of complex multiplications. In the following complexity analysis, for simplicity, we assume that $N,N_{t}\gg N_{r}$ which is a typical case for a SIM-based broadcast communication system.
@@ V-A DPC Precoding
The covariance matrix optimization is performed by Algorithm 1, with the complexity of obtaining $\mathbf{U}$ from $\mathbf{S}$ being negligible. In addition, the complexity of one iteration of inner loop (i.e., lines 2 to 5 in Algorithm 1) is $\mathcal{O}(N_{t}^{2}N_{r})$. Also, the complexity of line 6 is approximately $\mathcal{O}(KN_{t}N_{r}^{2})$, while the complexity of line 7 can be neglected. Let $I_{U}$ denote the number of inner loops (i.e., lines 1 to 1 in Algorithm 1), then the complexity of Algorithm 1 is given by $\mathcal{O}(I_{U}KN_{t}^{2}N_{r})$.
Optimizing the SIM phase shifts requires $\mathcal{O}(KNN_{t}N_{r})$ multiplications for computing $\sum\nolimits_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k}$. The matrix inversion $\mathbf{A}^{-1}$ and its multiplication contribute $\mathcal{O}(N_{t}^{3}+NN_{t}^{2})$ multiplications. Since $(\mathbf{W}^{m})^{H}(\boldsymbol{\mathbf{\Phi}}^{m})^{H}$ is precomputed and only the diagonal elements are needed in (35), the additional complexity is $\mathcal{O}(LN^{3})$. Hence, the complexity of the gradient calculation is $\mathcal{O}(NN_{t}^{2}+LN^{3})$. After obtaining $\boldsymbol{\phi}^{(n+1)}$, the complexity of calculating $\mathbf{B}$ and $\mathbf{H}$ is $\mathcal{O}(LN^{3})$. In addition, $\mathcal{O}(KN_{t}N_{r}(N_{t}+N_{r})+N_{t}^{3})\approx\mathcal{O}(N_{t}^{3})$ multiplications are needed to obtain $h(\boldsymbol{\phi}^{(n+1)})$. The complexity of the SIM phase shifts optimization is $\mathcal{O}(NN_{t}^{2}+I_{\phi,D}(LN^{3}+N_{t}^{3}))$, where $I_{\phi,D}$ is the number of line search iterations.
Therefore, the overall computational complexity for single iteration of Algorithm 2 is given by
$\!\!\!\!\!\!C_{\text{SIM-DPC}}$ $=\mathcal{O}(I_{U}KN_{t}^{2}N_{r}+NN_{t}^{2}+I_{\phi,D}(LN^{3}+N_{t}^{3})).$ (56)
V-B Linear Precoding (LP)
The optimization of the precoding matrices follows Algorithm 3, where computing $\bar{\mathbf{H}}$ requires $\mathcal{O}(K^{2}N_{t}N_{r}^{2})$ multiplications. Computing the initial $\mathbf{X}$ from $\mathbf{P}$ requires $\mathcal{O}(K^{2}N_{t}N_{r}^{2}+K^{3}N_{r}^{3})$ multiplications. Furthermore, the complexity of calculating all $\mathbf{\hat{Z}}_{k}$, $\mathbf{\hat{Y}}_{k}$ and $\mathbf{A}_{k}$ is $\mathcal{O}(K^{3}N_{r}^{3}+K^{2}N_{r}^{3})$. To determine $\mathbf{X}$,
  • if $N_{t}\geq KN_{r}$, solving (49) requires $\mathcal{O}(K^{3}N_{r}^{3})$ multiplications, or otherwise according to
  • if $N_{t}<KN_{r}$, solving (51) requires $\mathcal{O}(K^{3}N_{r}^{3}+K^{3}N_{t}N_{r}^{2})\approx\mathcal{O}(K^{3}N_{t}N_{r}^{2})$ multiplications.
Thus, the computational complexity of precoding matrix optimization is written as
@@ V-B Linear Precoding (LP)
The complexity for calculating $\sum_{k}\bar{R}_{k}$ and $f(\mathbf{X}^{(n)})$ can be neglected. Let $I_{X}$ denote the number of inner loops (i.e., lines 3 to 8 in Algorithm 3) , then the complexity of Algorithm 3 is given by $\mathcal{O}(I_{X}C_{x})$.
To optimize the phase shifts of the SIM, we need $\mathcal{O}(KN_{t}^{2})$ multiplications to obtain $\hat{\mathbf{P}}_{s}$ and all $\hat{\mathbf{P}}_{k}$ matrices, while calculating $\mathbf{F}_{1,k}$ and $\mathbf{F}_{2,k}$ requires $\mathcal{O}(KN_{t}N_{r}(N_{t}+N_{r}))$ multiplications, which also holds for $\mathbf{F}_{1,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{s}-\mathbf{F}_{2,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{k}$. Multiplying these terms with $\mathbf{G}_{k}^{H}$ adds $\mathcal{O}(KNN_{t}^{2})$ multiplications. As mentioned earlier, since $(\mathbf{W}^{m})^{H}(\boldsymbol{\mathbf{\Phi}}^{m})^{H}$ are precomputed and only the diagonal elements are need in (55), the additional complexity is $\mathcal{O}(LN^{3})$. Hence, the complexity of the gradient calculation is $\mathcal{O}(KNN_{t}^{2}+LN^{3})$. After obtaining $\boldsymbol{\phi}^{(n+1)}$, the complexity of calculating $\mathbf{B}$ and $\mathbf{H}$ is $\mathcal{O}(LN^{3})$. To obtain all terms $\sum\nolimits_{j}\mathbf{H}_{k}\mathbf{P}_{j}\mathbf{P}_{j}^{H}\mathbf{H}_{k}^{H},$ we need $\mathcal{O}(KN_{t}N_{r}(N_{t}+N_{r}))$ multiplications. Any additional complexity for computing $g(\boldsymbol{\phi}^{(n+1)})$ can be neglected. Hence, the complexity of optimizing the SIM phase shifts is $\mathcal{O}(KNN_{t}^{2}+I_{\phi,L}LN^{3})$, where $I_{\phi,L}$ is the number of line search steps.
Therefore, the overall computational complexity for one iteration of Algorithm 4 is given by
$C_{\text{SIM-LP}}=\mathcal{O}(I_{X}C_{x}+KNN_{t}^{2}+I_{\phi,L}LN^{3}).$ (58)